Can you help me please? Solve the system by the addition method.
x – y = 100
0.20x + 0.06y = 150
Thank you all

Answer by nerdybill(7384)

Starting with:
x – y = 100
0.20x + 0.06y = 150
.
Multiply both sides of the top equation by 0.06:
0.06x – 0.06y = 6
0.20x + 0.06y = 150
.
Now, add both top and bottom equation together:
0.06x – 0.06y = 6
0.20x + 0.06y = 150
----------------------
0.26x = 156
x = 156/0.26
x = 600
.
Substitute the above back into the first equation:
x – y = 100
600 – y = 100
-y = 400
y = -400
.
(x,y) = (600, -400)